OpenStudy (anonymous):

Linear Algebra Question The trace of a matrix is the sum of the entries along the main diagonal. Prove that if A is a 2 x 2 matrix with det(A) = 1, and trace(A) > 2, then A has two real and distinct eigenvalues. (answer please, will be offline until tomorrow)

OpenStudy (kinggeorge):

Well, if you write out your matrix as\[\left[\begin{matrix} a&b\\c&d \end{matrix}\right]\]Then we know that the eigenvalues are given by \[\left| \begin{matrix} (a-\lambda) &b\\c&(d-\lambda) \end{matrix} \right|=0\]so \[ad-\lambda (a+d)+\lambda^2 -bc = 0\]Now, notice that \(ad-bc\) is the determinant of A, and \(a+d\) is the trace of A. So we can rewrite this as \(\lambda^2-\lambda(a+d)+1=0\) Where \(a +d>2\). If we let \(a+d=2\), then we can factor the equation as \((\lambda -1)^2\) which only has 1 distinct root. Moreover, if we graph this, we see that the vertex touches the x-axis at \(x=1\). Now, note that if \(a+d>2\), then we are subtracting a larger number, so the graph of the function \(\lambda^2-\lambda(a+d)+1\) is shifted down. Since it is an upward opening parabola, that means that it must have two distinct roots if shifted down. QED